# SynthVLM: Towards High-Quality and Efficient Synthesis of Image-Caption Datasets for Vision-Language Models

### v4

@@ Abstract.
Vision-Language Models (VLMs) have recently emerged, demonstrating remarkable vision-understanding capabilities. However, training these models requires large-scale datasets, which brings challenges related to efficiency, effectiveness, quality, and privacy of web data. In this paper, we introduce SynthVLM, a novel data synthesis and curation method for generating image-caption pairs. Unlike traditional methods, where captions are generated from images, SynthVLM utilizes advanced diffusion models and high-quality captions to automatically synthesize and select high-resolution images from text descriptions, thereby creating precisely aligned image-text pairs. To demonstrate the power of SynthVLM, we introduce SynthVLM-100K, a high-quality dataset consisting of 100,000 curated and synthesized image-caption pairs. In both model and human evaluations, SynthVLM-100K outperforms traditional real-world datasets. Leveraging this dataset, we develop a new family of multimodal large language models (MLLMs), SynthVLM-7B and SynthVLM-13B, which achieve state-of-the-art (SOTA) performance on various vision question-answering (VQA) tasks. Notably, our models outperform LLaVA across most metrics with only 18% pretrain data. Furthermore, SynthVLM-7B and SynthVLM-13B attain SOTA performance on the MMLU benchmark, demonstrating that the high-quality SynthVLM-100K dataset preserves language abilities. To facilitate future research, our dataset and the complete data generating and curating methods are open-sourced at https://github.com/starriver030515/SynthVLM.
Synthetic Images, Data Curation, Vision Large Language Models
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: KDD ’25: ACM KDD conference; August 3–7, 2025; Toronto, ON, Canadabooktitle: KDD ’25: ACM KDD conference, August 3–7, 2025, Toronto, ON, Canadaisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
In recent years, with the rapid advancements in large language models (LLMs) (OpenAI, 2023a; Touvron et al., 2023) and multimodal large language models (MLLMs) (Zhao et al., 2023; Wu et al., 2023), data management has become a crucial aspect of these technologies (Fernandez et al., 2023; Trummer, 2023; Chen et al., 2023a; Miao et al., 2024; Nie et al., 2023; Huang et al., 2024). At the same time, Bai et al. (2024) also demonstrates that data processing, selection, and management can significantly influence the performance of MLLMs.
Among MLLMs, Vision Large Language Models (VLLMs) achieve competitive performance in traditional multimodal tasks such as image classification (Chen et al., 2024), image understanding (Li et al., 2023b, c), and image captioning (Bai et al., 2023). Most existing VLLMs rely on web images and primarily focus on modifying model architectures to better leverage information from multiple modalities (Li et al., 2023b, c; Bai et al., 2023; Liu et al., 2023b, a; Chen et al., 2024). However, high-quality web data is scarce, and the quality of data significantly impacts the performance of VLLMs. For example, Li et al. (2023d); Wang et al. (2024a) demonstrate that higher-quality training data can enhance the effectiveness of VLLMs. The key to ensuring high-quality data lies in the precise alignment between multimodal inputs, such as the alignment between captions and images.
To address the scarcity of high-quality web data, data synthesis strategies have increasingly been employed, particularly with the advancement of generative models. For instance, Nguyen et al. (2023) utilized BLIP2 to generate numerous image captions, achieving SOTA results on DataComp. In the domain of VLLMs, Chen et al. (2023b) used GPT-4 Vision to produce highly descriptive image captions, resulting in significant improvements in LLaVA. The integration of these generative models has opened new avenues for enhancing data quality and alignment, further boosting VLLM performance.
Despite these notable advancements in VLLMs, the challenge of multimodal data persists, particularly when paired image and text are required. Although synthetic captions have been generated for images, the absence of generated images means that the issue of limited high-quality web images remains unresolved. This limitation gives rise to four key challenges that still need to be addressed:
@@
Our evaluations using metrics such as CLIPScore  (Hessel et al., 2021) and training results reveal that existing datasets still align the modalities sub-optimally. Web images often suffer from issues such as blurriness and watermarks. Additionally, methods employing BLIP2  (Nguyen et al., 2023) for generating captions frequently result in logical inconsistencies and unclear grammar, which mislead VLLM training and degrade their language capabilities.
@@
Existing datasets frequently lack the necessary diversity and depth required to train VLLMs effectively. For example, many web-scraped datasets consist of low-quality or irrelevant content that fails to capture the complexity of natural language and visual information. Furthermore, the limited scope of current datasets often results in models that struggle to generalize across various contexts.
@@
Methods that rely on manual captioning are both labor-intensive and resource-intensive. Automated solutions such as ShareGPT4v (Chen et al., 2023b), which leverage GPT-4 Vision for labeling, are costly and difficult to scale. Additionally, many existing strategies require the creation of large datasets, such as LLAVA-558K, to boost performance, resulting in significant data redundancy.
C4. Security Risks.
To address these challenges, we introduced a new data generation pipeline, SynthVLM. First, we implemented a quality selection process to filter high-quality caption data. Next, we employed advanced diffusion models to generate images based on these captions. For quality assurance, we proposed a new method for evaluating and selecting image-caption pairs by combining CLIPScore (Hessel et al., 2021) and SSIM (Wang et al., 2004), focusing on both image-text alignment and image quality. Our data generation approach achieved superior alignment between images and captions compared to existing methods. With 100K curated synthetic data, we attained SOTA results across multiple benchmarks, utilizing only 18% of the official LLaVA-558K dataset size.
Overall, our contributions are as follows:
  • New Method. We propose a novel two-step image-caption data synthesis method called SynthVLM. Using SynthVLM, we create a high-quality image-caption dataset, SynthVLM-100K. We subsequently use SynthVLM-100K to train both the SynthVLM-7B and SynthVLM-13B models. Our method is both effective and efficient compared to the baseline method. Additionally, our ablation study demonstrates that both steps of SynthVLM are effective.
  • High Data Quality. SynthVLM-100K outperforms baseline datasets in both model and human evaluations. Specifically, SynthVLM-100K achieves superior CLIPScore and SSIMScore, indicating better image-text alignment and image quality. Moreover, it receives better ratings from human annotators, GPT-4, and Intern-VL, further confirming the dataset’s high-quality.
  • Strong Performance on VLLMs. We use only SynthVLM-100K as high-quality pre-training data, outperforms the baseline that uses LLaVA-558K. We achieve SOTA performance on vision-understanding tasks and demonstrate outstanding results on language tasks on MMLU benchmark.
  • Data Privacy. By using synthetic data, we eliminate the need for real personal images, documents, and other sensitive information, thereby ensuring data privacy.
@@ 2. Related Work
2.1. Diffusion Model
$q(x_{1},\ldots,x_{T}\mid x_{0}):=\prod_{t=1}^{T}q(x_{t}\mid x_{t-1}),$
where $\beta_{t}$ controls the variance of the noise added at each step.
The distribution after $t$ steps can be written as:
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}(1-\beta_{i})$.
While DDPMs have shown promising results, several improvements have been proposed to enhance their efficiency  (Song et al., 2021; Zhu et al., 2024) and sample quality  (Nichol and Dhariwal, 2021; Dhariwal and Nichol, 2021). The superior performance of diffusion models has been leveraged in various sub-tasks, including image generation, image translation, inpainting  (Lu et al., 2022; Su et al., 2023).
2.2. Vision Language Models
The integration of visual knowledge into large language models (LLMs) has become a pivotal area of research due to the rapid advancements in LLMs. VLLMs combine vision information from vision encoders with LLMs, thus enabling these models to process and interpret visual inputs for various visual tasks (Liu et al., 2023d; Zhang et al., 2022; Li et al., 2022b) with enhanced accuracy and efficiency. Pioneering frameworks like CLIP (Radford et al., 2021) leverage contrastive learning on expansive image-caption datasets to align modalities, forming the groundwork for cross-modal comprehension. Various adapters (Liu et al., 2023b, a; Li et al., 2023b, 2022a; Jian et al., 2023; Lu et al., 2023a) are introduced to further integrate different modalities. For example, LLaVA (Liu et al., 2023b, a) employs a straightforward MLP to inject the vision information into LLMs. Whereas more complex implementations like the Q-Former in BLIP (Li et al., 2022a, 2023b) utilize cross-attention to enhance modality integration.
Recent studies  (Wang et al., 2024b; Chen et al., 2023b; Liu et al., 2023b, a; Li et al., 2023f) aims to boost VLLM performance by focusing on the quality of both pre-training and fine-tuning datasets. Models like LLaVA  (Liu et al., 2023b, a) and ShareGPT4V  (Chen et al., 2023b) have shown remarkable advancements in understanding and following complex instructions through instruction tuning. Although these improvements help align the vision modality and establish a solid basis for cross-modal comprehension, they require extensive datasets for training and could potentially diminish the model’s language capabilities. In this work, we propose a new data generation strategy that leverages high-quality caption-image pairs, enabling the model to achieve state-of-the-art (SOTA) results with minimal data. By aligning during the pre-training (PT) phase, we have further enhanced the model’s language capabilities.
@@ 2.3. Data Quality and Selection
The advent of large language models has brought about a substantial increase in the volume of training data (Touvron et al., 2023; OpenAI, 2023b). In this scenario, the quality and quantity of data become paramount. LLMs, trained on vast amounts of data, can capture subtle nuances and complex patterns in language, excelling in various natural language processing tasks. However, the increase in data volume also brings new challenges, particularly in data management, cleaning, and annotation (Bai et al., 2024). In this section, we mainly discuss the effectiveness of data quality and data selection.
Data Quality
: High-quality data can significantly enhance the performance of models (meta llama, 2024). As the volume of data increases, ensuring high data quality becomes more challenging because it requires more resources for data cleaning, selection and annotation (Bai et al., 2024). Poor quality data can lead to models learning incorrect patterns and making inaccurate predictions.
Data Selection
: LLMs-based methods were commonly used in data selection (Bai et al., 2024). For instance, Du et al. (2023) leverages DeBERTa (He et al., 2020) for scoring, retaining high-quality data, and combining it with the k-center greedy algorithm to select diverse data. Chen et al. (2023c) score the accuracy of data using ChatGPT to pick out high-quality data. Xu et al. (2023b) use GPT-4 to rewrite data to increase their complexity and then streamline it by reducing its variety and improving its quality. Liu et al. (2023c) train two models using ChatGPT’s labeled data to score the quality and complexity of the data. Lu et al. (2023b) rely on ChatGPT to tag each instance, defining its complexity and diversity based on these tags. Parkar et al. (2024) first cluster the data, and then use GPT-4 to select high-quality data for each cluster.
Figure 2. We compared the SynthVLM-100K dataset with LLaVA-558K. As shown in (a), the generated image can avoid content such as watermarks and advertisements. In (b), the caption mentions a joicy named plate, however in the bottom right image, joicy is actually a book. The generated images better reflect the content of the captions. Additionally, the resolution of the generated images is 1024x1024, which is higher than the existing images, is more beneficial for model training and expansion.
@@ 2.4. Data Generation
Data has always been the key driver behind the success of LLMs. Recent advancements of LLMs largely due to the availability of large-scale, diverse, and high-quality datasets for training these models (Lin et al., 2023; Li et al., 2024c, e, 2023g, a, d, b). However, the scarcity of data and the high costs present substantial challenges in obtaining such datasets (Xie et al., 2024; Xu et al., 2023a). Recent advancements in generating synthetic data and improving the performance of LLMs have shown promising results across various domains. Synthetic data holds great potential in building large-scale, high-quality datasets. Researchers have explored multiple approaches, from leveraging differential privacy to creating instruction-tuning frameworks, to enhance the quality, diversity, and utility of synthetic data  (Wang et al., 2024c; Wei et al., 2023; Lou et al., 2023; Yang et al., 2023). A key component in generating high-quality synthetic datasets is precise alignment. Fan et al. (Fan et al., 2024) introduce REALIGN, a method that enhances the quality of instruction data by reformatting responses to better align with pre-established criteria and evidence, thereby improving LLMs’ alignment with human values while minimizing human annotation and model hallucinations. Li et al. (Li et al., 2023e) build a high-quality instruction-following language model by automatically labeling human-written text with corresponding instructions and demonstrating highly effective self-alignment.
@@ 3. Method
In this section, we introduce our data generation pipeline and then compare the generated dataset with other commonly used datasets. In subsection 3.1, we explore the construction of our caption set and the subsequent generation of 1M caption-image pairs using diffusion models. In subsection 3.2, we outline the filtering process applied to the generated 1 million image-text pairs. Then we meticulously select 100K high-quality, well-aligned caption-image pairs, ensuring the robustness and relevance of our dataset. In subsection 3.3, we demonstrate the effectiveness of our method by comparing the SynthVLM-100K dataset with other existing datasets.
@@ 3.1. Step1: Synthetic Dataset Construction
In this section, we introduce the image generation pipeline. First, we construct a large pool of captions for selection. Next, we generate high-quality images corresponding to these captions. We then select the best captions from the pool for image-text generation. Utilizing these high-quality captions, we employ diffusion models to generate the images.
@@
COCO-Caption COCO Human 118K
BLIP-LCS LCS BLIP 558K
ShareGPT4V LCS, COCO, etc GPT4-Vision 100K
ShareGPT4V-PT LCS, COCO, etc Share-Captioner 1246K
@@
Data Source.
To ensure the diversity of the captions, we combined human-generated and model-generated captions. As shown in Table  1. The human-generated captions were primarily sourced from LAION, CC, and SBU, while the model-generated captions were predominantly created using the method described in  (Nguyen et al., 2023), which utilizes BLIP2 to regenerate captions for images in the DataComp dataset  (Gadre et al., 2023).
Caption Curation.
To maintain dataset quality, we first removed low-quality captions, such as advertisements, overly repetitive descriptions, and captions with significant grammatical errors. The filtering process was performed with ChatGPT, combined with statistical indicators such as N-grams and Perplexity, ensuring that only high-quality, informative captions were used for training. For the remaining captions, we calculated the CLIPScore for these captions and their corresponding raw images. CLIPScore is a metric that measures the cosine similarity between images and their corresponding captions. The formula for calculating CLIPScore is as follows:
(1) ${CLIPScore}(I,C)=\frac{{CLIP}(I)\cdot{CLIP}(C)}{||{CLIP}(I)||\cdot||{CLIP}(C)||}$
where $I$ represents the image, $C$ represents the caption, and ${CLIP}(I)$ and ${CLIP}(C)$ denote the image and text feature vectors extracted by the CLIP model. The dot product of the vectors is denoted by $\cdot$, and |||||$\cdot$||| | ⋅ | | denotes the norm of the vectors.
@@
After filtering 1M high-quality captions, we employed Stable Diffusion XL (SDXL)  (Podell et al., 2023), a SOTA model capable of efficiently generating high-quality, high-resolution images. We configured SDXL with 60 sampling steps and used 8 A100 GPUs to ensure the superior quality of our synthetic images. Additionally, SynthVLM produces images at a resolution of 1024x1024, effectively addressing the low-resolution issues present in existing datasets. This improvement greatly enhances the quality and utility of the training data across various image generation and recognition tasks.
3.2. Step2: Synthetic Data Selection
In this section, we introduce a novel algorithm for quality control of generated datasets. To better ensure the alignment between images and their corresponding text descriptions, we continue to use CLIPScore. For a given image $I$ and its corresponding text $C$, we first calculate the image-text CLIPScore, ${CLIPScore}(I,C)$, to assess their alignment. Additionally, the generated images have a resolution of $1024\times 1024$, which will be resized to $336\times 336$ for compatibility with CLIP. To account for the potential loss in quality due to this resizing process, we introduce the Structural Similarity Index Measure (SSIM), a metric for image quality control. This will help us better ensure the image quality is preserved after resizing. The SSIM formula is defined as follows:
$I_{{resized}}={resize}({resize}(I,(336,336)),(1024,1024))$
The SSIMScore is computed as:
$SSIMScore=SSIM(I,I_{{resized}})$
Sample Model Sythetic Image win Web Image win
1K GPT4-Vision 633 367
1K InternVL2 692 308
1K Human 758 242
Datasets # Samples CLIPScore SSIMScore Weighted_Score
COCO-Caption 118K 0.31 0.73 0.67
BLIP-LCS 558K 0.32 0.75 0.70
ShareGPT4V 100K 0.32 0.79 0.71
Synth Dataset 1000K 0.34 0.78 0.73
Curated-Synth Dataset 100K 0.36 0.86 0.79
Methods SynthVLM LLaVA w/o selection
Dataset Number (k) 100 558 1000
Data Usage 33MB 27GB 330MB
@@ 3.2. Step2: Synthetic Data Selection
Figure 3. This is our process and prompt design for match assessment using GPT4-Vision. We consider various aspects, including the quality of the image and the match between the image and the caption, to help GPT4-Vision make a better selection. Based on this process, we compare SynthVLM with existing datasets from the model’s perspective.

### v5

@@ Abstract.
Vision-Language Models (VLMs) have recently emerged, demonstrating remarkable vision-understanding capabilities. However, training these models requires large-scale datasets, which brings challenges related to efficiency, effectiveness, and quality of web data. In this paper, we introduce SynthVLM, a new data synthesis and curation method for generating image-caption pairs. Unlike traditional methods, where captions are generated from images, SynthVLM utilizes advanced diffusion models and high-quality captions to synthesize and select images from text captions, thereby creating precisely aligned image-text pairs. We further introduce SynthVLM-100K, a high-quality dataset consisting of 100K curated and synthesized image-caption pairs. In both model and human evaluations, SynthVLM-100K outperforms traditional real-world datasets. Leveraging this dataset, we develop a new family of multimodal large language models (MLLMs), SynthVLM-7B and SynthVLM-13B, which achieve state-of-the-art (SOTA) performance on various vision question-answering (VQA) tasks. Notably, our models outperform LLaVA across most metrics with only 18% pretrain data. Furthermore, SynthVLM-7B and SynthVLM-13B attain SOTA performance on the MMLU benchmark, demonstrating that the high-quality SynthVLM-100K dataset preserves language abilities. Our dataset and the complete data generating and curating methods can be found in https://github.com/starriver030515/SynthVLM.
Synthetic Data, Data Curation, Vision Language Models
[…]
$\dagger$ Corresponding Author
Figure 1. We compared SynthVLM-100K with LLaVA-558K. In (a), generated image can avoid content such as watermarks and advertisements. In (b), the generated images better reflect the content of the captions. Additionally, the resolution of the generated images is higher than real images.
@@ 1. Introduction
In recent years, with the rapid advancements in large language models (LLMs) (OpenAI, 2023; Touvron et al., 2023) and multimodal large language models (MLLMs) (Zhao et al., 2023; Wu et al., 2023), data management has become a crucial aspect of these technologies (Fernandez et al., 2023; Trummer, 2023; Chen et al., 2023a; Miao et al., 2024; Nie et al., 2023; Huang et al., 2024). At the same time,  (Bai et al., 2024; Zhu et al., 2025; Zhou et al., 2024; Zhu et al., 2023) also demonstrates that data processing, selection, and management can significantly influence the performance of MLLMs.
Vision Language Models (VLMs), a subset of MLLMs, excel in tasks like image classification, understanding, and captioning (Chen et al., 2024; Li et al., 2023c, d; Bai et al., 2023). While most VLMs focus on improving model architectures to integrate multimodal data (Li et al., 2023c, d; Bai et al., 2023; Liu et al., 2023b, a; Chen et al., 2024), their reliance on web-sourced data poses a bottleneck: high-quality, precisely aligned image-text pairs are scarce. Noisy or misaligned data directly limits performance, as shown by  (Li et al., 2023e; Wang et al., 2024a), who link enhanced VLM effectiveness to carefully curated datasets. Crucially, precise alignment between modalities underpins data quality, which is a challenge demanding greater focus than architectural refinements alone.
To address the scarcity of high-quality web data, data synthesis strategies have increasingly been employed (Gao et al., 2025; Nguyen et al., 2023; Chen et al., 2023b). For instance,  (Nguyen et al., 2023) utilized BLIP2 to generate numerous image captions, achieving SOTA results on DataComp. In the domain of VLMs,  (Chen et al., 2023b) used GPT-4 Vision to produce highly descriptive image captions, resulting in significant improvements in LLaVA. The integration of these generative models has opened new avenues for enhancing data quality and alignment, further boosting VLM performance.
Despite these notable advancements in VLMs, the challenge of multimodal data persists, particularly when paired image and text are required. Although synthetic captions have been generated for images, the absence of generated images means that the issue of limited high-quality web images remains unresolved. This limitation gives rise to three key challenges that still need to be addressed:
@@
Web-sourced images often contain artifacts such as blurriness and watermarks, compromising quality. Furthermore, caption generation approaches (Nguyen et al., 2023) leveraging BLIP2 tend to introduce logical inconsistencies and grammatical errors in text, which propagate through VLM training pipelines, ultimately weakening their linguistic reasoning capabilities.
@@
Existing datasets frequently lack the diversity and depth required to train VLMs effectively. Many web-scraped datasets consist of low-quality or irrelevant content that fails to capture the complexity of natural language and visual information. Furthermore, the limited scope of current datasets often results in models that struggle to generalize across various contexts.
@@
Methods that rely on manual captioning are both labor-intensive and resource-intensive. Automated solutions such as ShareGPT4V (Chen et al., 2023b), which leverage GPT-4 for labeling, are costly and difficult to scale. Additionally, many strategies require the creation of large datasets, resulting in significant data redundancy.
To address these challenges, we introduced a new data generation pipeline, SynthVLM. First, we implemented a quality selection process to filter high-quality caption data. Next, we employed diffusion models to generate images based on these captions. For quality assurance, we proposed a new method for evaluating and selecting image-caption pairs by combining CLIPScore (Hessel et al., 2021) and SSIM (Wang et al., 2004), focusing on both image-text alignment and image quality. Our data generation approach achieved superior alignment between images and captions compared to existing methods. With 100K curated synthetic data, we attained SOTA results across multiple benchmarks, utilizing only 18% of the official LLaVA-558K dataset size. Overall, our contributions are as follows:
  • Innovative Data Synthesis Framework. We introduce SynthVLM, a novel two-stage pipeline for generating synthetic image-caption pairs, and curate SynthVLM-100K—a large-scale, high-quality dataset produced through this framework. Compared to conventional approaches, SynthVLM achieves superior efficiency without compromising quality.
  • Superior Synthetic Data Quality. SynthVLM-100K outperforms baseline datasets in both model and human evaluations. It achieves superior CLIPScore and SSIMScore, while also earning higher human evaluation ratings. Further validation via GPT-4 and Intern-VL2 assessments corroborates its enhanced quality for training vision-language models.
  • State-of-the-Art Vision-Language Modeling. When pretrained exclusively on SynthVLM-100K, our models surpass baseline methods relying on LLaVA-558K, achieving top-tier performance on vision-language understanding benchmarks. SynthVLM-13B also delivers exceptional performance on the MMLU benchmark, highlighting its multimodal versatility.
@@ 2.1. Data Quality and Selection
High-quality data can significantly enhance the performance of models (meta llama, 2024). Ensuring high data quality becomes more challenging because it requires more resources for data cleaning, selection and annotation (Bai et al., 2024). LLMs-based methods were commonly used in data selection (Bai et al., 2024). For instance,  (Du et al., 2023) leverages DeBERTa (He et al., 2020) for scoring, retaining high-quality data, and combining it with the k-center greedy algorithm to select diverse data.  (Chen et al., 2023c) scores the accuracy of data using ChatGPT to pick out high-quality data.  (Xu et al., 2023) use GPT-4 to rewrite data to increase their complexity and then streamline it by reducing its variety and improving its quality.
@@ 2.2. Data Generation
Recent advancements in generating synthetic data and improving the performance of LLMs have shown promising results across various domains. A key component in generating high-quality synthetic datasets is precise alignment.  (Fan et al., 2024) introduce REALIGN, a method that enhances the quality of instruction data by reformatting responses to better align with pre-established criteria and evidence, thereby improving LLMs’ alignment with human values while minimizing human annotation and model hallucinations.  (Li et al., 2023f) build a high-quality instruction-following language model by automatically labeling human-written text with corresponding instructions and demonstrating highly effective self-alignment.
@@ 3.1. Step1: Synthetic Dataset Construction
In this section, we introduce the image generation pipeline. First, we construct a large pool of captions. We then select the best captions from the pool for image-text generation. Utilizing these high-quality captions, we employ diffusion models to generate the images.
Data Source.
To ensure the diversity of the captions, we combined human-generated and model-generated captions. As shown in Table  1. The human-generated captions were primarily sourced from LAION, CC, and SBU, while the model-generated captions were created using the method described in  (Nguyen et al., 2023), which utilizes BLIP2 to regenerate captions for images in the DataComp dataset  (Gadre et al., 2023).
Caption Curation.
To maintain dataset quality, we first removed low-quality captions, such as advertisements, overly repetitive descriptions, and captions with significant grammatical errors. The filtering process was performed with ChatGPT, combined with statistical indicators such as N-grams and Perplexity, ensuring that only high-quality, informative captions were used for training.
@@
COCO-Caption (Lin et al., 2015) COCO Human 118K
BLIP-LCS (Liu et al., 2023a) LCS BLIP 558K
ShareGPT4V (Chen et al., 2023b) LCS, COCO, etc GPT4-Vision 100K
ShareGPT4V-PT (Chen et al., 2023b) LCS, COCO, etc Share-Captioner 1246K
@@
For the remaining captions, we calculated the CLIPScore(Hessel et al., 2021) for these captions and their corresponding raw images. CLIPScore is a metric that measures the cosine similarity between images and their corresponding captions. The formula for calculating CLIPScore is as follows:
[…]
where $I$ represents the image, $C$ represents the caption, and CLIP(I){CLIP}(I) and CLIP(C){CLIP}(C) denote the image and text feature vectors extracted by the CLIP model. The dot product of the vectors is denoted by $\cdot$, and $||\cdot||$ denotes the norm of the vectors.
@@
After filtering 1M high-quality captions, we employed Stable Diffusion XL (SDXL)  (Podell et al., 2023), a SOTA model capable of efficiently generating high-quality, high-resolution images. SynthVLM produces images at a resolution of 1024x1024, effectively addressing the low-resolution issues present in existing datasets. This improvement greatly enhances the quality and utility of the training data across various image generation and recognition tasks.
@@
3.2. Step2: Synthetic Data Selection
In this section, we introduce a novel algorithm for quality control of generated datasets. To better ensure the alignment between images and their corresponding text descriptions, we continue to use CLIPScore. For a given image $I$ and its corresponding text $C$, we first calculate the image-text CLIPScore, CLIPScore(I,C){CLIPScore}(I,C), to assess their alignment. Additionally, the generated images have a resolution of $1024\times 1024$, which will be resized to $336\times 336$ for compatibility with CLIP (Radford et al., 2021). To account for the potential loss in quality due to this resizing process, we introduce the Structural Similarity Index Measure (SSIM), a metric for image quality control. This will help us better ensure the image quality is preserved after resizing. The SSIM formula is defined as follows:
$I_{{resized}}={resize}({resize}(I,(336,336)),(1024,1024))$
The SSIMScore is computed as:
$SSIMScore=SSIM(I,I_{{resized}})$
Datasets # Samples CLIPScore SSIMScore Weighted_Score
COCO-Caption 118K 0.31 0.73 0.67
BLIP-LCS 558K 0.32 0.75 0.70
ShareGPT4V 100K 0.32 0.79 0.71
Synth Dataset 1000K 0.34 0.78 0.73
Curated-Synth Dataset 100K 0.36 0.86 0.79
Figure 3. Our process and prompt design for match assessment using GPT4V. We consider various aspects, including the quality of the image and the match between the image and the caption. Based on this process, we compare SynthVLM with existing datasets from the model’s perspective.